OpenStudy (anonymous):

PLEASE HELP WITH GRAPH AND THREE PART QUESTION!!!!!!! I HAVE BUMPED IT 3 TIMES ALREADY A recent poll yielded the following: (L) (h) (S) Total Male(M): 600 365 535 1500 Female(F): 418 559 523 1500 Total: 1018 924 1058 3000 Find: Probability(M union H) Probability(L') Probability(F intersect S)

OpenStudy (anonymous):

P(M Union H) = (1500 + 559) / 3000 = 2059/300 there were 1500 males in the the survey, and 559 Females supported H P(L ') = (3000 - 1018) / 3000 = 1982 / 3000 = 991/ 1500 P (Female intersect S) = 523 / 3000 There were 523 Females who supported S, that is the intersection